# Interpreting Equivariant Representations

### v1

@@ 3.2 When no isometric cross section exists: Random invariant linear projection
Random projections are often available:  Ma et al. (2018); Maron et al. (2019) propose a basis for all permutation invariant linear maps, and (Cesa et al., 2022a) describe how $E(n)$-equivariant and invariant linear maps can be constructed from a basis. Initializing these layers at random, we obtain an analogy to the random projections known from classical statistics (Candes & Tao, 2006).
As we cannot generally invert random projection, these invariant representations do not allow interpolation-based analysis. However, they are still highly valuable for visualization and interpretation, as well as building new models and analyses directly from the invariant latent representation.
4 Relation to existing invariant, quotient and equivariant latent spaces
Having presented our methods and notation, we can now explain in detail how they relate to recent related work. While the utility and interpretability of equivariant latent spaces is, to the best of our knowledge, unexpored in the past, the autoencoder literature explores latent space design. Mehr et al. (2018) design a quotient autoencoder for 3D shapes whose latents reside on the quotient space $\mathcal{X}/G$. Here, $\mathcal{X}$ parametrizes 3D shape and $G$ is the group of rotations or non-rigid deformations. As the latent space is $G$-invariant, shape alignment and interpolation are greatly simplified. However, as discussed above, quotient spaces are often non-Euclidean, greatly hindering their applicability.
Graph autoencoders commonly use a permutation invariant encoder. This can be achieved by using permutation equivariant layers (e.g. graph-convolutions) eventually followed by a permutation invariant layer Winter et al. (2021). As the composition of equivariant and invariant functions is invariant, this ensures that the latent representation is indeed invariant to permutation of the input nodes. This was the strategy followed e.g. in early graph VAE models Simonovsky & Komodakis (2018); Vignac & Frossard (2021); Rigoni et al. (2020); Liu et al. (2018), where an expensive graph alignment step was needed in order to train the model. To counteract this, newer models Hy & Kondor (2023) replace the invariant latent space with equivariant ones, similar to those studied in this paper. Similarly to the early VAE models, Winter et al. (2022) construct encoder-decoder architectures; here, however, the needed alignment of outputs with inputs is learned rather than optimized.
@@ 4 Relation to existing invariant, quotient and equivariant latent spaces
[…]
More generally, any invariant latent feature embedding $f_{inv}\colon\mathcal{X}\to\mathcal{Z}_{inv}$ which carries enough information to enable decoding back onto $\mathcal{X}$, as in (Winter et al., 2022), can necessarily be written as a composition $f_{proj}\circ f_{quot}$ of a quotient latent feature embedding $f_{quot}\colon\mathcal{X}\to\mathcal{Z}/G$ for some latent space $\mathcal{Z}$, and a projection-like map $f_{proj}\colon\mathcal{Z}/G\to\mathcal{Z}_{inv}$. Thus, any differences in performance as observed in the experiments of Winter et al. (2022), are caused by implementation choices rather than differences in the underlying mathematical model – invariant, quotient and equivariant representations are, mathematically, able to carry the same information. We argue, that when taking care to respect the group action when utilizing and interpreting the equivariant latent representation $\mathcal{Z}$, there is no good reason to avoid it.
A final existing approach to obtaining invariant and equivariant maps is using fundamental domains Aslan et al. (2023). When isometric cross sections, they form fundamental domains. However, you cannot generally find isometric cross sections – indeed, the quotient space can exhibit strongly non-Euclidean geometry Calissano et al. (2023). This explains, in part, why the quotient itself is often not an efficient invariant representation. This also indicates that making an isometric, or even near-isometric, mapping to the representation space mathematically impossible. Random invariant projections, on the other hand, are in our cases designed as continuous mappings from the equivariant feature space, and they therefore preserve some local geometric structure.
@@ 5 Experiments
5.1 Isometric invariant representations via a graph VAE
We consider a permutation equivariant graph variational autoencoder (VAE) $h\colon\mathcal{X}\to\mathcal{Z}\to\mathcal{X}$ trained for molecule generation, and evaluate the utility of its latent codes for various downstream tasks. Each graph $(V,E)\in\mathcal{X}$ consists of a set $V$ of (at most) $n$ nodes with an $n\times d_{A}$ node feature matrix, and an $n\times n\times d_{E}$ edge feature tensor $E$, where $n$ denotes the number of nodes in the graph. A permutation $g\in S_{n}$ acts on the graph $(V,E)$ through its associated permutation matrix $P_{g}$:
@@ 5.1 Isometric invariant representations via a graph VAE
Figure 2: The first two principal components of the training data for the equivariant (top) and invariant representation (bottom). Each column illustrates a specific molecular property.
Dataset.
The QM9 dataset Ramakrishnan et al. (2014); Ruddigkeit et al. (2012) consists of approx. 130.000 stable, small molecules, using 80%/10%/10% for training/validation/testing. Each molecule is represented by at most 9 heavy atoms, their bindings (edges) and selected molecular properties. As our interest is in the permutation equivariant representation, we simplify the graphs to contain only the atom-type (node features) and binding-type (edge features). All graphs are padded with not-a-node and not-an-edge features to obtain the same node number for each graph.
Model.
Our permutation equivariant VAE is based on linear equivariant layers as derived in Maron et al. (2018b), combined with entry-wise non-linearities, which were used in both the encoder and decoder. A comprehensive description of the model architecture can be found in appendix.
Visualisation.
@@
The top row of Fig. 2 shows the first two principal components of the equivariant latent representation [^1] Note that the equivariant representations were obtained by randomly permuting the input nodes to remove any implicit ordering which may have been implied by the structure of the dataset. $\mathcal{Z}$ of the QM9 test set, with color representing different molecular properties. Inspecting the equivariant latent representations in the top row incorrectly suggests no apparent structure in the data, as molecules with similar molecular properties are by no means close in the latent space. On the other hand, when inspecting the isometric invariant representation $\mathcal{Z}_{s}$ plotted in the bottom row, a different picture emerges. Here, we find a clear pattern between latent representation and molecular properties, indicating that the model does indeed pick up on important structure in the data. In other words, looking at the same latent representation using either the equivariant representation, or its invariant projection, leads to very different conclusions about both the data and the model.
@@
Figure 4: We investigate the pairwise Hamming distance between consecutive observation pairs sampled along the linear interpolation. Right: Histograms over the average pairwise distance for $10000$ interpolations in the equivariant (blue) and invariant (red) representations. Left: visualization of the pairwise hamming distance along the single interpolation shown in Fig. 3.
Fig. 3 compares the equivariant interpolation $s(\alpha z_{i}+(1-\alpha)z_{j})$ to the isometric, invariant interpolation $\alpha s(z_{i})+(1-\alpha)s(z_{j})$ by visualizing 25 decoded molecules sampled along each of the interpolating lines (left). The same two interpolating lines are compared through the lens of the isometric, invariant representation $\mathcal{Z}_{s}$ (right). We see that interpolation in the latent space yields pathological behavior, where molecule structure varies greatly along the line, whereas interpolation using the isometric, invariant cross section yields a far more smooth variation along the curve.
We confirm this tendency of difference in interpolation smoothness by computing the pairwise Hamming distance between consecutive molecules sampled along the interpolating line. Fig. 4 shows quantitatively that consecutive graphs generated by interpolating in the equivariant representation, are less stable than those generated by interpolating in the invariant, isometric representation.
Stability of molecules in a neighbourhood.
Using the test-set containing approximately 13.000 molecules, i.e. 10% of the available data, we now compare local structure of the equivariant- and invariant representations using a $k$-nearest neighbour regression model to predict molecular properties from latent representations. We report the mean absolute error (MAE), as is the custom for QM9 Wu et al. (2018), as a function of the number of neighbours considered, for each of the properties in figure 5. In addition to the isometric invariant representation, we also consider several invariant pooling operations commonly used to summarize equivariant representations. It is clear that the MAE is consistently lowest when training a predictor on the isometric invariant representations, indicating that they are better at preserving information needed for downstream analysis.
@@ 5.2 Rotation invariant MNIST classifier
Figure 6: F1 score of KNN classifier.
@@
The first two principal components of the equivariant and invariant representations are shown in Fig. 7. We see that naively using the equivariant representation yields a plot suggesting very little signal. However, after applying a random linear projection, the structure becomes strikingly clear. We analyze the local structure of the representations by applying a $k$-nearest neighbour classifier. Its quality is evaluated using the F1 score; see Fig. 6. Again, we see that the quality of the classifier increases when applied to the invariant representation as opposed to the equivariant one.
Figure 7: The figure shows the first two principal components of the training data for the equivariant representation (right), and the invariant representation (left). The colours represent the labels.
@@ 6 Discussion
Weaknesses & future work
@@
Invariant and equivariant models rely on advanced mathematical machinery that makes them less accessible than standard deep learning models. One might be tempted to take our results as an argument to just avoid these models altogether. However, this would not suffice to avoid our illustrated problems. An alternative and common way to encourage invariance or equivariance is to rely on the inherent flexibility of deep learning models to learn approximate invariance/equivariance through data augmentation. Fig. 8 illustrates the effect of using augmentation to encourage rotation invariance in a classification CNN trained on MNIST Deng (2012) – we see a very similar pattern as in the previously described $SO(2)$ invariant classifier (right). On the left, we see how an intermediate latent representation of the CNN represents the MNIST test images when they are rotated randomly (”Augmented Dataset”, left) or embedded using their original orientation (”Aligned Dataset”, right). Since MNIST images have an orientation, they are naturally aligned with each other, and the resulting t-SNE plot nicely separates the different digits, obtaining an effect similar to the $SO(2)$ invariant classifier. For most image classification problems, however, there is no natural alignment, and a more realistic scenario is the plot on the left, where random rotations have been applied prior to embedding. What this shows is that while a pre-embedding alignment can work as a poor man’s invariant representation, the lack of well-defined relative distances that we saw for strictly equivariant representations cannot be avoided by sticking to more straightforward models.
Figure 8: For the latent representation of a CNN classifier trained with rotation augmentation, rotated MNIST images (left) exhibit a similar embedding behavior as the equivariant representation of a $SO(2)$ invariant classifier, while the image alignment given by the natural orientation of digits provides a poor man’s analogue of the invariant representation for the $SO(2)$ invariant classifier.
@@ A.1 Proof of proposition 1
A.2 Proof of proposition 2
@@ Appendix B Permutation equivariant VAE
By Maron et al. (2018a; 2019); Thiede et al. (2020); Pan & Kondor (2022) we have that we can define any linear function $L:\mathbb{R}^{n^{k}\times d}\to\mathbb{R}^{n^{l}\times d^{\prime}}$ using exactly $b(k+l)dd^{\prime}$ known basis elements, where $b(\cdot)$ denotes the Bell number. Using this result, we can define node- and edge-level linear equivariant layers:
@@
by using the a weighted linear combination of the known basis elements. That amounts to a total of $8d_{v}d^{\prime}_{v}$ weights in the case of $L_{V}$ and $15d_{e}d^{\prime}_{e}$ weights in the case of $L_{E}$. For an exact construction $L_{V}$ and $L_{E}$ please refer to Pan & Kondor (2022). Using this construction we can define a linear layer $L_{V,E}:\mathbb{R}^{n\times d_{v}}\times\mathbb{R}^{n^{2}\times d_{E}}\to\mathbb{R}^{n^{2}\times d^{\prime}_{E}}$, by the channel-wise concatenation of $L(V)$ and $L(E)$. Note, that this concatenation does not change the equivariance property of the layer. All linear layers of the architecture utilised in the current work is of one of these forms. Subsequently a ReLU activation function is applied.
@@
Each of the linear layers, except for the last layer, are followed by 2D convolutions, ReLU activations and instance normalization as described above.
@@
  • The last layer is then the concatenation of a linear layer mapping between matrix representations (the reconstructed edge-matrix), and a linear layer mapping to a feature vector representation (the reconstructed node-matrix). The reconstructed edgematrix is enforced to be symmetric by adding the transpose of itself.
@@
The model was trained using the negative evidence lower bound (ELBO) as is standard for VAEs. A learning rate of 0.0001 and a batch-size of 32 was chosen. The model was trained for 1000 epochs. The QM9 dataset was obtained through the pytorch-geometric library Fey & Lenssen (2019).

### v2

@@ 4.2 A Choice of Continuous Map: Random Invariant Linear Projections
Random projections are often available:  (Ma et al., 2018; Maron et al., 2019) propose a basis for all permutation invariant linear maps, and (Cesa et al., 2022a) describe how $E(n)$-equivariant and invariant linear maps can be constructed from a basis. Initializing these layers at random, we obtain an analogy to the random projections known from classical statistics (Candes & Tao, 2006).
As we cannot generally invert invariant random projections, they do not allow interpolation-based analysis. However, they are still highly valuable for visualization and interpretation, as well as building new models and analyses directly from the invariant latent representation.
5 Relation to Existing Invariant-, Quotient- and Equivariant Latent Spaces
Having presented our methods and notation, we can now explain in detail how they relate to recent related work. While the utility and interpretability of equivariant latent spaces is, to the best of our knowledge, unexplored in the past, the autoencoder literature explores latent space design. (Mehr et al., 2018) design a quotient autoencoder for 3D shapes whose latents reside on the quotient space $\mathcal{X}/G$. Here, $\mathcal{X}$ parametrizes 3D shape and $G$ is the group of rotations or non-rigid deformations. As the latent space is $G$-invariant, shape alignment and interpolation are greatly simplified. However, as discussed above, quotient spaces are often non-Euclidean, greatly hindering their applicability.
Graph autoencoders commonly use a permutation invariant encoder. This can be achieved by using permutation equivariant layers (e.g. graph-convolutions) eventually followed by a permutation invariant layer (Winter et al., 2021). As the composition of equivariant and invariant functions is invariant, this ensures that the latent representation is indeed invariant to permutation of the input nodes. This was the strategy followed e.g. in early graph VAE models (Simonovsky & Komodakis, 2018; Vignac & Frossard, 2021; Rigoni et al., 2020; Liu et al., 2018), where an expensive graph alignment step was needed in order to train the model. To counteract this, newer models (Hy & Kondor, 2023) replace the invariant latent space with equivariant ones, similar to those studied in this paper. Similarly to the early VAE models, (Winter et al., 2022) construct encoder-decoder architectures; here, however, the needed alignment of outputs with inputs is learned rather than optimized.
@@ 5 Relation to Existing Invariant-, Quotient- and Equivariant Latent Spaces
${\mathcal{X}}$${\mathcal{Z}}$${\mathcal{Y}}$${\mathcal{Z}/G}$${\mathcal{Z_{\textit{inv}}}}$$\scriptstyle{f}$$\scriptstyle{f_{quot}}$$\scriptstyle{g}$$\scriptstyle{\pi}$$\scriptstyle{f_{proj}}$
Figure 4: The first two principal components of the QM9 training dataset for the equivariant (top) and invariant representation (bottom) of the latent space. Each column illustrates a specific molecular property.
More generally, any invariant latent feature embedding $f_{inv}\colon\mathcal{X}\to\mathcal{Z}_{inv}$ which carries enough information to enable decoding back onto $\mathcal{X}$, as in (Winter et al., 2022), can necessarily be written as a composition $f_{proj}\circ f_{quot}$ of a quotient latent feature embedding $f_{quot}\colon\mathcal{X}\to\mathcal{Z}/G$ for some latent space $\mathcal{Z}$, and a projection-like map $f_{proj}\colon\mathcal{Z}/G\to\mathcal{Z}_{inv}$. Thus, any differences in performance as observed in the experiments of (Winter et al., 2022), are caused by implementation choices rather than differences in the underlying mathematical model – invariant, quotient and equivariant representations are, mathematically, able to carry the same information. We argue that when taking care to respect the group action when utilizing and interpreting the equivariant latent representation $\mathcal{Z}$, there is no good reason to avoid it.
Another existing approach to obtaining invariant and equivariant maps is using fundamental domains (Aslan et al., 2023). When isometric cross sections, they form fundamental domains. However, in general, one cannot find isometric cross sections – indeed, the quotient space can exhibit strongly non-Euclidean geometry (Calissano et al., 2024). This explains, in part, why the quotient itself is often not an efficient invariant representation. This also indicates that making an isometric, or even near-isometric, mapping to the representation space mathematically impossible. Random invariant projections, on the other hand, are in our cases designed as continuous mappings from the equivariant feature space and, therefore, preserve some local geometric structure.
@@ 6 Experiments
6.1 Isometric Invariant Representations via Equivariant Grap VAE
We consider a permutation equivariant graph variational autoencoder (VAE) $h\colon\mathcal{X}\to\mathcal{Z}\to\mathcal{X}$ trained for molecule generation, and evaluate the utility of its latent codes for downstream tasks. A graph $(V,E)\in\mathcal{X}$ consists of a set $V$ of (at most) $n$ nodes with an $n\times d_{A}$ node feature matrix and an $n\times n\times d_{E}$ edge feature tensor $E$, where $n$ denotes the number of nodes. A permutation $g\in S_{n}$ acts on the graph $(V,E)$ through its associated permutation matrix $P_{g}$:
@@ 6.1 Isometric Invariant Representations via Equivariant Grap VAE
Dataset:
The QM9 dataset (Ramakrishnan et al., 2014; Ruddigkeit et al., 2012) consists of approx. 130.000 stable, small molecules, using 80%/10%/10% for training/validation/testing. Each molecule is represented by at most 9 heavy atoms, their bindings (edges), and selected molecular properties. As our interest is in the permutation equivariant representation, we simplify the graphs to contain only atom-type (node features) and binding-type (edge features). Each graphs is padded with not-a-node and not-an-edge features to obtain the same number of nodes.
Model:
Our permutation equivariant VAE is based on linear equivariant layers as derived in (Maron et al., 2018b), combined with entry-wise nonlinearities, which were used in both the encoder and decoder. A comprehensive description of the model architecture can be found in Appendix B.
Visualisation:
@@
The upper row of Fig. 4 shows the first two principal components of the equivariant latent representation [^1] Note that the equivariant representations were obtained by randomly permuting the input nodes to remove any implicit ordering which may have been implied by the structure of the dataset. $\mathcal{Z}$ of the QM9 test set, with molecular properties encoded by color. Inspecting the equivariant latent representations in the top row incorrectly suggests no apparent structure in the data, as molecules with similar molecular properties are by no means close in the latent space. On the other hand, when inspecting the isometric invariant representation $\mathcal{Z}_{s}$ plotted in the bottom row, a different picture emerges. Here, we find a clear pattern between latent representation and molecular properties, indicating that the model does indeed pick up on important structure in the data. In other words, looking at the same latent representation using either the equivariant representation, or its invariant projection, leads to very different conclusions about both the data and the model.
@@
Figure 6: We measure pairwise Hamming distances between consecutive observation pairs sampled along linear interpolations. Right: Histograms over the average pairwise distance for $10000$ interpolations in the equivariant (blue) and invariant (red) representations. Left: visualization of the pairwise Hamming distance along the single interpolation from Fig. 5.
Fig. 5 compares the equivariant interpolation $s(\alpha z_{i}+(1-\alpha)z_{j})$ with the isometric invariant interpolation $\alpha s(z_{i})+(1-\alpha)s(z_{j})$ by visualizing 25 decoded molecules sampled along each of the interpolating lines (left). The same two interpolating lines are compared through the lens of the isometric, invariant representation $\mathcal{Z}_{s}$ (right). We see that interpolation in the latent space yields pathological behavior, where molecule structure varies greatly along the line, whereas interpolation using the isometric, invariant cross section yields a far more smooth variation along the curve.
@@
We confirm this tendency of difference in interpolation smoothness by computing the pairwise Hamming distance between consecutive molecules sampled along the interpolating line. Fig. 6 shows quantitatively that consecutive graphs generated by interpolating in the equivariant representation are less stable than those generated by interpolating in the invariant, isometric representation.
Stability of Molecules in a Neighborhood.
Using the test-set containing approximately 13.000 molecules, that is, 10% of the available data, we now compare the local structure of the equivariant and invariant representations using a $k$-nearest neighbor regression model to predict molecular properties from latent representations. We report the mean absolute error (MAE), as is custom for QM9 (Wu et al., 2018), as a function of the number of neighbors considered, for each of the properties in Figure 7. In addition to the isometric invariant representation, we also consider several invariant pooling operations commonly used to summarize equivariant representations. It is clear that the MAE is consistently lowest when training a predictor on the isometric invariant representations, indicating that they are better at preserving information needed for downstream analysis.
@@ 6.2 Rotation Invariant MNIST Classifier
Figure 8: MNIST dataset classification problem: F1 score of KNN classifier applied to Invariant and Equivariant representations of the latent space.
@@
The first two principal components of the equivariant and invariant representations are shown in Fig. 9. We see that naively using the equivariant representation yields a plot suggesting very little signal. However, after applying a random linear projection, the structure becomes strikingly clear. We analyze the local structure of the representations by applying a $k$-nearest neighbor classifier. Its quality is evaluated using the F1 score; see Fig. 8. Again, we see that the quality of the classifier increases when applied to the invariant representation as opposed to the equivariant one.
Figure 9: The figure shows the first two principal components of the MNIST training data for the equivariant representation (right), and the invariant representation (left) of the latent space. The colours represent the classifier labels.
@@ 7 Discussion
Figure 10: t-SNE plots showing the demonstrating how image alignment given by the natural orientation of digits provides a poor man’s analogue of the invariant representation for the $SO(2)$ invariant classifier
Weaknesses & future work:
@@
Invariant and equivariant models rely on advanced mathematical machinery that makes them less accessible than standard deep learning models. One might be tempted to take our results as an argument to just avoid these models altogether. However, this would not suffice to avoid our illustrated problems. An alternative and common way to encourage invariance or equivariance is to rely on the inherent flexibility of deep learning models to learn approximate invariance/equivariance through data augmentation. Fig. 10 illustrates the effect of using augmentation to encourage rotation invariance in a classification CNN trained on MNIST (Deng, 2012) – we see a very similar pattern as in the previously described $SO(2)$ invariant classifier (bottom). On the top plot, we see how an intermediate latent representation of the CNN represents the MNIST test images when they are rotated randomly (”Augmented Dataset”, left) or embedded using their original orientation (”Aligned Dataset”, right). Since MNIST images have an orientation, they are naturally aligned with each other, and the resulting t-SNE plot nicely separates the different digits, obtaining an effect similar to the $SO(2)$ invariant classifier. For most image classification problems, however, there is no natural alignment, and a more realistic scenario is the plot on the left, where random rotations have been applied prior to embedding. What this shows is that while a pre-embedding alignment can work as a poor man’s invariant representation, the lack of well-defined relative distances that we saw for strictly equivariant representations cannot be avoided by sticking to more straightforward models.
Acknowledgements
This work was supported in part by the Independent Research Fund Denmark (grant no. 1032-00349B), by the Novo Nordisk Foundation through the Center for Basic Machine Learning Research in Life Science (grant no. NNF20OC0062606), the Pioneer Centre for AI, DNRF grant nr P1, and by the ERC Advanced grant $786854$ on Geometric Statistics.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Proofs
A.2 Proof of Proposition 4.4
The following proof is heavily inspired by the proof of Theorem 22.2. and Corollary 22.3 from (Munkres, 2000), but is included for the sake of completeness.
We first show that $s^{\prime}$ is continuous if and only if $s$ is continuous. First assume $s$ is continuous. As the canonical projection $\pi:\mathcal{Z}\to\mathcal{Z}/G$ is a quotient map, hence continuous, the composition $s^{\prime}\circ\pi=s$ is continuous. Now suppose $s$ is continuous. Then given an open set $U\subseteq\mathcal{Z}_{s}$ we have that $s^{-1}(U)$ is open in $\mathcal{Z}$. But since the diagram of proposition 4.1 commutes, we have $s^{-1}(U)=(s^{\prime}\circ\pi)^{-1}(U)$. Since $\pi$ is a quotient it now follows that $s^{\prime-1}(U)$ is open in $\mathcal{Z}/G$.
Next, we show, that $s^{\prime}$ is a quotient map if and only if $s$ is a quotient map. Assume $s^{\prime}$ is a quotient map. Again, the composition $s^{\prime}\circ\pi=s$ is a quotient map, as the composition of two quotient maps is itself a quotient map. Now suppose that $s$ is a quotient map. Since $s$ is surjective, we have that $s^{\prime}\circ\pi=s$ is surjective, and hence $s^{\prime}$ is surjective. Now consider a subset $U\subseteq\mathcal{Z}_{s}$, and assume $s^{\prime-1}(U)$ is open in $\mathcal{Z}/G$. Then $\pi^{-1}(s^{\prime-1}(U))$ is open in $\mathcal{Z}$ as $\pi$ is continuous. But then $s^{(}-1)(U)=\pi^{-1}(s^{\prime-1}(U))$ is also open in $\mathcal{Z}$, and as such, $U$ is open in $\mathcal{Z}_{s}$ as $s$ is a quotient map. In combination with the fact that $s^{\prime}$ is continuous if and only if $s$ is continuous the assertion follows.
Lastly the assertion that $s^{\prime}$ is a homeomorphism if and only if $s^{\prime}$ is a bijective quotient map follows directly from the definitions.
A.3 Proof of Proposition 4.5
@@ Appendix B Permutation equivariant VAE
By (Maron et al., 2018a, 2019; Thiede et al., 2020; Pan & Kondor, 2022) we have that we can define any linear function $L:\mathbb{R}^{n^{k}\times d}\to\mathbb{R}^{n^{l}\times d^{\prime}}$ using exactly $b(k+l)dd^{\prime}$ known basis elements, where $b(\cdot)$ denotes the Bell number. Using this result, we can define node- and edge-level linear equivariant layers:
@@
by using the a weighted linear combination of the known basis elements. This amounts to a total of $8d_{v}d^{\prime}_{v}$ weights in the case of $L_{V}$ and $15d_{e}d^{\prime}_{e}$ weights in the case of $L_{E}$. For an exact construction $L_{V}$ and $L_{E}$ please refer to (Pan & Kondor, 2022). Using this construction, we can define a linear layer $L_{V,E}:\mathbb{R}^{n\times d_{v}}\times\mathbb{R}^{n^{2}\times d_{E}}\to\mathbb{R}^{n^{2}\times d^{\prime}_{E}}$, by the channel-wise concatenation of $L(V)$ and $L(E)$. Note that this concatenation does not change the equivariance property of the layer. All linear layers of the architecture utilised in the current work is of one of these forms. Subsequently, a ReLU activation function is applied.
@@
Each of the linear layers, except for the last layer, are followed by 2D convolutions, ReLU activations, and instance normalization as described above.
@@
  • The last layer is then the concatenation of a linear layer mapping between matrix representations (the reconstructed edge-matrix) and a linear layer mapping to a feature vector representation (the reconstructed node-matrix). The reconstructed edgematrix is enforced to be symmetric by adding the transpose of itself.
@@
The model was trained using the negative evidence lower bound (ELBO) as is standard for VAEs. A learning rate of 0.0001 and a batch-size of 32 was chosen. The model was trained for 1000 epochs. The QM9 dataset was obtained through the Python Geometric library (Fey & Lenssen, 2019).
B.1 Model
We assume the following likelihood model of the examined graphs:
$\log p_{\theta}(G|\mathbf{z})=\log p_{\theta}(V|\mathbf{z})+\log p_{\theta}(E|\mathbf{z})$ (18)
From the perspective of a variational autoencoder, $\theta$ denotes the parameters of the decoder. We assume that each node vector $\mathbf{v}\in V$ and edge vector $\mathbf{c}\in E$ are obtained independently from a categorical distribution parameterized by $\theta$. That is:
$p_{\theta}(V|\mathbf{z})$ $=\prod_{\mathbf{v}\in V}p_{\theta}(\mathbf{v}|\mathbf{z})$ (19)
$p_{\theta}(E|\mathbf{z})$ $=\prod_{\mathbf{c}\in E}p_{\theta}(\mathbf{c}|\mathbf{z}),$ (20)
where $p_{\theta}(\mathbf{v}|\mathbf{z})$ and $p_{\theta}(\mathbf{c}|\mathbf{z})$ respectfully denotes the probablility of observing $\mathbf{v}$ and $\mathbf{c}$ respectfully under the model.
Following the standard notation for VAEs, we impose the prior $p(\mathbf{z})=\mathcal{N}(\textbf{0}|\mathbf{1})$, and as a consequence we naturally let the approximate posterior be given as:
$q_{\phi}(\mathbf{z}|G)=\mathcal{N}(\mathbf{\mu}_{\phi}(G)|\mathbf{\sigma}^{2}_{\mathbf{\phi}}(G))$ (21)
where, $\phi$ denotes the parameters of the encoder of the VAE. Our objective is to minimize the negative evidence lower bound (ELBO) of the log-likelihood $\log(p_{\theta}(G))$, i.e.:
$-ELBO$ $=\mathbb{E}_{\mathbf{z}\sim q_{\phi}(\mathbf{z}|G)}[-\log p_{\theta}(G|\mathbf{z}))]+KL(q_{\phi}(\mathbf{z}|G)||p_{\theta}(\mathbf{z}))$ (22)
$=\mathbb{E}_{\mathbf{z}\sim q_{\phi}(\mathbf{z}|G)}\left[-\log\left(\frac{p_{\theta}(G|\mathbf{z})p(\mathbf{z})}{q_{\phi}(\mathbf{z}|G)}\right)\right],$ (23)
where $KL(\cdot||\cdot)$ denotes the Kullback Leibler (KL) divergence.